# Quasi-likelihood inference for SDE with mixed-effects observed at high frequency

### v1

@@ A.3.1. Stochastic expansions of $\widehat{M}_{i}$ and $\widehat{v}_{i}$
$|\check{\mathsf{R}}^{M}_{i}|\leq\sqrt{h}\,|\overline{\mathfrak{m}}_{i}|\left(1+\frac{1}{\sqrt{nN}}|\widehat{u}_{\eta}|\right).$ (A.82)
Thus, applying (A.2) in Lemma A.1 gives
$\widehat{M}_{i}=\tau_{i}\int_{0}^{T}S_{i}(t)^{-1}a_{i}(t)^{\otimes 2}dt+\mathsf{R}^{M}_{i}=:M_{i,T}+\mathsf{R}^{M}_{i}$ (A.83)
@@
In a similar manner to [6, Lemma 2], we can show that
@@
for any Riemann-integrable $\mathsf{f}(y)$ for which $\sup_{i}\sup_{t\leq T}E[|\mathsf{f}(Y_{i}(t))|^{K}]<\infty$ for every $K>0$. This implies that $v_{1,i}=\overline{\mathfrak{m}}_{i}$. It is easy to see that
$v_{2,i}=\frac{\sqrt{h}}{nN}\sum_{j}\overline{\mathfrak{b}}_{ij}a_{i,j-1}y_{ij}=\frac{\sqrt{h}}{N}\,\overline{\mathfrak{m}}_{i}.$ (A.87)
@@
A.3.2. Estimating the gap between $\mathbb{H}_{2}(\vartheta_{2})$ and $\mathbb{H}^{\star}_{2}(\vartheta_{2})$
From the expressions (3.17) and (A.80), some algebra gives the estimate
@@
By (2.11), we can find a constant $\mathsf{c}_{0}>0$ such that $N/n^{1-\mathsf{c}_{0}}\to 0$. By (A.28),
@@
for any (small) $\mathsf{c}>0$, hence in particular
$\sqrt{N}\max_{i\leq N}|\widehat{M}_{i}-M_{i}|=o_{p}(\sqrt{N}h^{1/2-\mathsf{c}})=o_{p}(1).$ (A.93)
This implies that $P[B_{n}]\to 0$ where
$B_{n}:=\left\{\max_{i\leq N}\left|\frac{\widehat{M}_{i}-M_{i}}{M_{i}}\right|>h^{(1-\mathsf{c}_{0})/2}\right\}.$ (A.94)
We then obtain for any $Q>0$ and sufficiently large $n$,
$P[\mathcal{G}_{1,1}>Q]\leq P[B_{n}]+P\left[B_{n}^{c}\cap\{\mathcal{G}_{1,1}>Q\}\right]\to 0,$ (A.95)
concluding that $\mathcal{G}_{1,1}=o_{p}(1)$. Similarly, we can prove that $\mathcal{G}_{1,2}=p_{(}1)$ (much easier since $\Sigma\widehat{M}_{i}+I_{p_{\varphi}}\geq I_{p_{\varphi}}>O$). As for $\mathcal{G}_{2}$, some manipulations and estimates together with (A.93) give
$\mathcal{G}_{2}$ $\lesssim o_{p}(1)\frac{1}{N}\sum_{i}|\overline{\mathfrak{m}}_{i}||M_{i}^{-1}||\widehat{M}_{i}^{-1}|(1+|M_{i}^{-1}|).$ (A.96)
Assumption 3.5 ensures that
$\max\left\{\frac{1}{N}\sum_{i}\lambda_{\min}(M_{i})^{-2-\varepsilon_{1}},\leavevmode\nobreak\ \frac{1}{N}\sum_{i}\lambda_{\min}(\widehat{M}_{i})^{-1-\varepsilon_{2}}\right\}=O_{p}(1).$ (A.97)
Hence, by means of the generalized Hölder inequality, the upper bound in (A.96) turns out to be $o_{p}(1)$. Thus, it follows that
@@
A.3.3. Asymptotics of $\widehat{\vartheta}_{2}$
Building on what we have seen in the previous section, we can follow the $M$-estimation machinery for the i.i.d. model to obtain the consistency and the asymptotic normality of $\widehat{\vartheta}_{2}$. To derive the form of the asymptotic covariance matrix, we can follow the same line as in [5, Section 6.2].

### v2

@@ A.3.1. Stochastic expansions of $\widehat{M}_{i}$ and $\widehat{v}_{i}$
$|\check{\mathsf{R}}^{M}_{i}|\leq\sqrt{h}\,|\overline{\mathfrak{m}}_{i}|\left(1+\frac{1}{\sqrt{N}}|\widehat{u}_{\eta}|\right).$ (A.84)
Then, applying (A.4) in Lemma A.1 gives
$\widehat{M}_{i}=\tau_{i}\int_{0}^{T}S_{i}(t)^{-1}a_{i}(t)^{\otimes 2}dt+\mathsf{R}^{M}_{i}=M_{i,T}+\mathsf{R}^{M}_{i}$ (A.85)
@@
In a similar manner to [delattre2018parametric, Lemma 2], we can show that
@@
for any measurable $\mathsf{f}(y)$ satisfying $\sup_{i}\sup_{t\leq T}E[|\mathsf{f}(Y_{i}(t))|^{K}]<\infty$ for every $K>0$. This implies that $v_{1,i}=\overline{\mathfrak{m}}_{i}$. It is easy to see that
$v_{2,i}=\frac{\sqrt{h}}{nN}\sum_{j}\overline{\mathfrak{b}}_{ij}a_{i,j-1}\overline{y}_{ij}=\frac{\sqrt{h}}{N}\,\overline{\mathfrak{m}}_{i}.$ (A.89)
@@
In sum, we obtained
$|\widehat{M}_{i}-M_{i}|\leq\sqrt{h}\,|\overline{\mathfrak{m}}_{i}|\left(1+\frac{1}{\sqrt{N}}|\widehat{u}_{\eta}|\right),$ (A.92)
$|\widehat{v}_{i}-v_{i}|\leq\sqrt{h}\,|\overline{\mathfrak{m}}_{i}|\left(1+\frac{1}{\sqrt{nN}}|\widehat{u}_{\eta}|+\frac{\sqrt{h}}{N}|\widehat{u}_{\eta}|^{2}\right).$ (A.93)
A.3.2. Estimating the gap
From the expressions (3.17) and (A.82), some calculations give the estimate
@@ A.3.2. Estimating the gap
To deduce that all three terms on the rightmost side are negligible, we assume without loss of generality that the variables are scalar.
@@
By (2.11), we have $N/n^{\mathsf{a}^{\prime\prime}}\lesssim 1$ ($\mathsf{a}^{\prime\prime}\in(0,1)$). By (A.30) and (A.92),
@@
for any (small) $\mathsf{c}>0$, hence in particular taking $\mathsf{c}=(1-\mathsf{a}^{\prime\prime})/2$ gives
$\sqrt{N}\max_{i\leq N}|\widehat{M}_{i}-M_{i}|=o_{p}\big(\sqrt{N}h^{1/2-\mathsf{c}}\big)=o_{p}\big(N/n^{\mathsf{a}^{\prime\prime}}\big)=o_{p}(1).$ (A.97)
This and Assumption 3.5 imply that $P[B_{n}]\to 0$ where
$B_{n}:=\left\{\max_{i\leq N}\left|\frac{\widehat{M}_{i}-M_{i}}{M_{i}}\right|>h^{\mathsf{a}^{\prime\prime}/2}\right\}.$ (A.98)
For any $\mathsf{Q}>0$ and sufficiently large $n$,
$P[\mathcal{G}_{1,1}>\mathsf{Q}]\leq P[B_{n}]+P\left[B_{n}^{c}\cap\{\mathcal{G}_{1,1}>\mathsf{Q}\}\right]\to 0,$ (A.99)
concluding that $\mathcal{G}_{1,1}=o_{p}(1)$. Similarly, we can prove that $\mathcal{G}_{1,2}=o_{p}(1)$ (much easier since $\Sigma\widehat{M}_{i}+I_{p_{\varphi}}\geq I_{p_{\varphi}}>O$). As for $\mathcal{G}_{2}$, under Assumption 3.5, some manipulations and estimates together with (A.97) give
$\mathcal{G}_{2}$ $\lesssim o_{p}(1)\,\frac{1}{N}\sum_{i}|\overline{\mathfrak{m}}_{i}||M_{i}^{-1}||\widehat{M}_{i}^{-1}|(1+|M_{i}^{-1}|)$
$\leq o_{p}(1)\,\frac{1}{N}\sum_{i}|\overline{\mathfrak{m}}_{i}|=o_{p}(1).$ (A.100)
Thus, it follows that
@@
Analogously, we can also deduce that
$\sup_{\vartheta_{2}}\left|\frac{1}{\sqrt{N}}\left(\partial_{\vartheta_{2}}^{k}\mathbb{H}_{2}(\vartheta_{2})-\partial_{\vartheta_{2}}^{k}\mathbb{H}^{\star}_{2}(\vartheta_{2})\right)\right|=o_{p}(1)$ (A.102)
for $k=1,2,3$.
A.3.3. Asymptotic normality of $\widehat{\vartheta}_{2}$
Now we are in the position to conclude the consistency and the asymptotic normality of $\widehat{\vartheta}_{2}$. Building on (A.101) and (A.102), under Assumption 3.5 we can follow the standard $M$-estimation machinery for the i.i.d. model associated with the random function $\mathbb{H}^{\star}_{2}(\vartheta_{2})$ (see, for example, [vdV98, Chapter 5]). Finally, to derive the form of the asymptotic covariance matrix, we can follow the same line as in [DelGenLar18-2, Section 6.2].
Acknowledgement
We are grateful to the INRAE MIGALE bioinformatics platform https://migale.inrae.fr for providing computational resources.
This work was partially supported by JST CREST Grant Number JPMJCR2115, Japan, and by JSPS KAKENHI Grant Numbers 23K22410 and 24K21516 (HM).